OpenStudy (anonymous):

Find the equation of the line that passes through the point (6,9) and is parallel to y = -4/3x + 9 a. y=-4/3x +17 b. y=-4/3x + 1

OpenStudy (anonymous):

do you know what the slope is in y = -4/3x + 9 ??

OpenStudy (anonymous):

Yes, -4/3. :)

OpenStudy (anonymous):

good..now a parallel line will have the same slope, -4/3. Now we can use y = mx + b and solve for b. y = mx + b (6,9) x = 6 and y = 9 m = -4/3 now we sub 9 = -4/3(6) + b 9 = -24/3 + b 9 = -8 + b 9 + 8 = b 17 = b so your parallel equation is : y = -4/3x + 17

OpenStudy (anonymous):

Thank you so much!! xx

OpenStudy (anonymous):

no problem :)